Session: Nov 1999 (session 4)
Wavelength: 6 cm
To give feedback on this experiment, click
Medicina (A.Scalambra at Wed Nov 24 18:20:13 MET 1999) Minor failures
To beginning no correct tracking, just a few second.
Torun (K.Borkowski at Thu Nov 25 09:46:58 MET 1999) Minor failures
The experiment went without any problems and I would
normally classify it as a success. However yeaterday
Dave Graham informed us that Torun station again has
a mix up of tracks similar to that from the September
session. So it seems all the data from this session
till this experiment will be only marginally useful.
Noto (C.Trigilio at Thu Nov 25 11:06:53 MET 1999) Success
Effelsberg (A. Kraus at Thu Nov 25 11:18:54 MET 1999) Success
Westerbork (Tony Foley at Thu Nov 25 12:08:30 MET 1999) Minor failures
From 14:40UT one of the 13 telescopes in the array was mispointing.
Otherwise no known problems. Tape playback looks good.
Jodrell (Alastair Gunn at Wed Dec 1 03:39:27 MET 1999) Success
New receiver installed prior to this experiment due to cryogenics
failure overnight. Phase cal switched off (too strong). No known
Onsala (Johnn Conway at Mon Dec 13 09:54:53 MET 1999) Success
No knowbn problems
Calibration (Jean\-Francois Desmurs at Wed Dec 15 13:17:07 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER
in the nov99 directory under the name:
Read the comments in the file to get further information.